OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8 −96 0 24 48

OpenStudy (anonymous):

i will medal

zepdrix (zepdrix):

Hey William :)\[\large\rm \color{orangered}{f(x)=6x+12},\qquad\color{royalblue}{g(x)=x-8}\]

zepdrix (zepdrix):

Hmm so we need to solve this:\[\large\rm \color{orangered}{f\left[\color{royalblue}{g(6)}\right]}\]

zepdrix (zepdrix):

We want to work from the inside and move outward. So let's deal with the blue first. What is the value of the function g, when 6 is plugged in for x?

OpenStudy (anonymous):

so you say to turn it into g(x)=6-8?

zepdrix (zepdrix):

Good, yes.\[\large\rm g(x)=x-8,\qquad g(6)=6-8\]Looks like we end up with -2, ya?

zepdrix (zepdrix):

So then,

zepdrix (zepdrix):

That simplifies down the blue inner part for us. \[\large\rm \color{orangered}{f\left[\color{royalblue}{g(6)}\right]}=\color{orangered}{f\left[\color{royalblue}{-2}\right]}\]

zepdrix (zepdrix):

So now all we need to do is simply plug -2 into our f function.

OpenStudy (anonymous):

im having trouble someone talk me through this

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

is it f(x)=6(-2) +12

jimthompson5910 (jim_thompson5910):

yes now evaluate 6(-2) + 12 and you're done

OpenStudy (jdoe0001):

hmmm what would you get for say \(\bf {\color{brown}{ g(6)}}?\)

OpenStudy (anonymous):

0

OpenStudy (jdoe0001):

0? how.. did you get 0 though?

OpenStudy (jdoe0001):

anyhow \(\bf g(x)=x-8\qquad g({\color{brown}{ 6}})={\color{brown}{ 6}}-8\)

OpenStudy (anonymous):